OpenStudy (kimberly101):

⇒x+(x+5)=31 Solve X.

OpenStudy (anonymous):

Use the associative property of addition and combine like terms, then subtract five from both sides, then divide by 2 to solve.

OpenStudy (kimberly101):

ok thanks!!!

OpenStudy (anonymous):

Let me know what you get as a result :)

OpenStudy (kimberly101):

x is 13

OpenStudy (anonymous):

That is correct :D You can know that it is correct for certain by plugging 13 into the original equation for x, and you see that 13+(13+5)=13+18=31. Good job!

OpenStudy (kimberly101):

thanks!!!